### v1

# GAL-MAD: Towards Explainable Anomaly Detection in Microservice Applications Using Graph Attention Networks

Abstract
The transition to microservices has revolutionized software architectures, offering enhanced scalability and modularity. However, the distributed and dynamic nature of microservices introduces complexities in ensuring system reliability, making anomaly detection crucial for maintaining performance and functionality. Anomalies stemming from network and performance issues must be swiftly identified and addressed. Existing anomaly detection techniques often rely on statistical models or machine learning methods that struggle with the high-dimensional, interdependent data inherent in microservice applications. Current techniques and available datasets predominantly focus on system traces and logs, limiting their ability to support advanced detection models. This paper addresses these gaps by introducing the RS-Anomic dataset generated using the open-source RobotShop microservice application. The dataset captures multivariate performance metrics and response times under normal and anomalous conditions, encompassing ten types of anomalies. We propose a novel anomaly detection model called Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD), leveraging Graph Attention and Long Short-Term Memory architectures to capture spatial and temporal dependencies in microservices. We utilize SHAP values to localize anomalous services and identify root causes to enhance explainability. Experimental results demonstrate that GAL-MAD outperforms state-of-the-art models on the RS-Anomic dataset, achieving higher accuracy and recall across varying anomaly rates. The explanations provide actionable insights into service anomalies, which benefits system administrators.
Keywords Anomaly detection  $\cdot$ Microservices $\cdot$ Multivariate data  $\cdot$ Time series data  $\cdot$ Graph Attention  $\cdot$ Explainability
1 Introduction
The migration from monolithic to microservice architecture has become a prevalent trend in cloud applications due to benefits such as improved scalability, modularity, and autonomous components that simplify development. A few enterprise companies using microservice architecture include Amazon, Netflix, and Uber. A microservice application consists of loosely coupled, independently deployable services, with each microservice performing a specific functionality or a business objective. Typically the performance of microservices are gaurded by Service Level Obejctives (SLOs).
When a system experiences a fault or is being used unexpectedly, it can result in erroneous behaviour, also known as an anomaly. Detecting these anomalies is essential to ensure the entire microservice application functions correctly and efficiently. Several factors can contribute to anomalous behaviour, such as hardware faults, misconfigurations, and unexpected load. However, identifying and resolving these anomalies can be challenging due to the complexity of the microservices architecture, which involves multiple independent services working together, asynchronous communication, frequent changes, and complex dependencies. Even a minor change in one service can significantly impact others, making it difficult to isolate the root cause of the anomaly.
The existing datasets for anomaly detection in microservice applications primarily consist of system traces and logs, and most of them are not publicly available Zhao et al. (2020); Wu et al. (2020); Kohyarnejadfard et al. (2022). Further, many of these datasets contain only univariate time-series data, which lack sufficient detail and do not capture the interdependencies among services. As a result, they can lead to suboptimal performance in anomaly detection Zhao et al. (2020). Instead, a publicly available multivariate time-series dataset that includes microservice application performance data would significantly enhance anomaly detection in such applications.
Additionally, current approaches to anomaly detection in microservices primarily rely on statistical and machine learning techniques that focus on univariate or log-based data. While these methods can be effective in specific situations, they often overlook the complex interdependencies and high-dimensional characteristics of microservices. The accuracy and efficiency of these anomaly detection methods can vary significantly depending on the application domain and the type of data collected. Moreover, existing methods typically do not provide insights into the root causes of detected anomalies. Without a clear understanding of these root causes, system administrators face challenges in implementing timely and effective resolutions.
In this research, we present the RS-Anomic dataset, which contains multivariate data from a microservice application and a framework for detecting anomalous behaviour of a microservice application and localizing the affected microservice. We utilized RobotShop Instana (2022) as the microservice application for this dataset. RobotShop is a small-scale microservice application that aims to teach containerized application orchestration and monitoring methods. RS-Anomic provides performance metrics on memory, CPU utilization, file Input/Output, and network as a multivariate time series for each service. These metrics can be valuable in developing novel anomaly detection methods that explore inter-dependencies between services and their performance metrics.
Moreover, we introduce a new model called Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) for detecting anomalies in microservices. This model is an encoder-decoder architecture trained using the performance metrics of the microservice system during normal operation. When provided with the performance metrics over a specific period, the model learns to reconstruct this input accurately. During inference, if the system’s performance metrics indicate anomalous behaviour, the model exhibits a higher reconstruction loss, aiding in identifying such anomalies. The encoder component employs Graph Attention layers to capture the dependencies among cooperating microservices, and Long Short-Term Memory layers to recognize the temporal patterns in the operations of the microservice system. The decoder mirrors the structure of the encoder. To assess the performance of our model, we compare it with state-of-the-art anomaly detection models using the RS-Anomic dataset, highlighting the importance of capturing interdependencies within the microservices domain. Additionally, we utilize SHapley Additive exPlanations Lundberg and Lee (2017a) for root cause analysis, which reinforces the model’s predictions and enhances its practical utility.
2 Related Work
This section presents the existing work related to our research under three categories: datasets for anomaly detection, anomaly detection methods, and root cause analysis.
2.1 Datasets for Anomaly Detection
Performance and response time features are crucial in detecting and localizing anomalies in microservices. Large datasets are available for anomaly detection using microservice trace logs Qiu et al. (2020); Huye et al. (2023); Lee et al. (2024). By analyzing trace logs, slow or erroneous services can be identified. Traces serve as a unified response time metric for service calls, encompassing both latency and performance aspects. However, root cause analysis using traces are often limited to identifying the specific service call and the anomalous service. To analyze system level resource usage multivariate data consisting of performance metrics are required. Response time is a key component of latency related SLOs and is essential for anomaly detection. NAB dataset collection from Numeta Lavin and Ahmad (2015) consists of a dataset built using AWS server metrics collected by the Amazon CloudWatch service. However, this dataset lacks response time related metrics. In contrast, a more comprehensive dataset combining system resource utilization and response time was introduced by Luo et al. Luo et al. (2021), however it does not include network resource usage data. Since detecting fine network anomalies requires detailed network resource usage metrics, future datasets should incorporate these to enhance anomaly detection capabilities.
2.2 Anomaly Detection
In early anomaly detection techniques, statistical methods were commonly used. Researchers analyzed the probability distribution of data points, identifying points outside the defined range of normal behaviour as outliers, and therefore, anomalies. Traditional techniques mainly include non-deep learning and log analysis techniques. Log analysis is complicated as logs are unstructured for large systems and challenging to perform in real time. In addition, to log analysis techniques, traditional machine learning methods have been used alongside deep learning approaches.
Recent research has shown that deep learning-based approaches generally outperform traditional and statistical methods in anomaly detection. However, anomaly detection remains a difficult task due to the lack of anomalous data and the highly imbalanced nature of available datasets, as anomalies are rare compared to normal behaviour. To overcome this issue, unsupervised approaches were utilized to distinguish abnormal from normal patterns Liu et al. (2020); Nguyen et al. (2021); Chen et al. (2017).
There are several advanced models for detecting anomalies, including the Graph Deviation Network (GDN) Deng and Hooi (2021), the Multivariate Anomaly Detection with Generative Adversarial Network (MAD-GAN)Li et al. (2019), and KitsuneMirsky et al. (2018). In Chen et al. (2017) introduced autoencoder ensembles for unsupervised outlier detection, demonstrating that neural network-based anomaly detection techniques yield high accuracy. Their model reconstructs normal data, identifying anomalies as instances that cannot be accurately reconstructed. The problem of accurately learning complex dependencies in multivariate data is addressed by Graph Attention Networks (GAT) Velickovic et al. (2017). GAT is further extended in GDN Deng and Hooi (2021), using a cosine similarity to create a graph structure to represent the pairwise relationship between data points. It then uses graph attention-based forecasting to predict future values and calculates the graph deviation score to evaluate its performance. MAD-GAN Li et al. (2019) is designed to detect anomalies in time series data produced by Cyber-Physical Systems. This model is trained using normal data and uses a Long-Short Term Memory-Recurrent Neural Network discriminator and a reconstruction-based method to calculate the anomaly score for each sample. Kitsune Mirsky et al. (2018) is an unsupervised Network Intrusion Detection System that can detect local network attacks without supervision. Kitsune uses a collection of variational autoencoders trained to replicate network traffic patterns and gradually improve their performance with Kitsune’s core algorithm, KitNET, through incremental training and efficient pattern tracking.
2.3 Root Cause Analysis
There is a lack of research exploring the use of explainability in root cause analysis for anomaly detection. By employing a graph-based anomaly detection method that preserves the microservice architecture, we investigate the potential of anomaly localization through explainability. We propose locating anomalous services using explainability techniques Lundberg and Lee (2017b) in conjunction with the output of the model. The tension between accuracy and interpretability in complex models has led to the development of methods for interpreting predictions. Lundberg et. al. introduced SHAP, a unified framework that assigns importance values to input features in a prediction. This framework offers a unique solution with desirable properties, enhancing existing methods to provide improved alignment with human intuition.
3 Methodology
3.1 RS-Anomic Dataset
The RS-Anomic dataset comprises performance metrics for 12 services, each with 19 performance metrics and a variable number of response time metrics per service, as detailed in Table 1. Furthermore, the anomaly data in the RS-Anomic dataset covers ten anomalous behaviours that may occur in microservice applications. RS-Anomic contains 100464 normal and 14112 anomalous instances. Each microservice communicates with a different number of microservices, and the response times for each communication link are recorded in the dataset. The dataset and data loading scripts are available at https://github.com/ms-anomaly/rs-anomic. Normal data and anomalous data are contained in two zip files. Each zip file contains 2 folders named cAdvisor and response_times containing performance metrics and response time data respectively. Response time and performance data should be concatenated to obtain the complete dataset. In the case of anomalous data, cAdvisor and response_times folders are further divided into each anomaly type, to obtain the complete anomalous data, response times of each anomaly should be concatenated with the corresponding performance metrics.
3.1.1 Testbed and Data Collection Environment
The RS-Anomic dataset was created using RobotShop Instana (2022), a microservice application that implements containerized orchestration and monitoring techniques. RobotShop is an open-source project with an e-commerce web application built using a microservices architecture. All services are deployed as Docker containers on a single server, communicating through a Docker bridge network. The application comprises 12 services, with their dependencies illustrated in Figure 1. This application was used in Harlicaj (2021) to evaluate their model. The testbed was run on a server with a 40-core CPU and 64 GB Memory running Ubuntu 20.04.5. We used Prometheus Rabenstein and Volz (2015) for data acquisition. The Prometheus server was configured to poll data from the microservices every 5 seconds. We instrumented the response times in microservice calls using Prometheus client libraries. The instrumented services are highlighted in Figure 1. Using Prometheus client libraries response time metrics were exposed from the microservices over HTTP endpoints. The cumulative summation of response time and the number of service calls that occurred between data polling for each communication link were recorded. Container Advisor (cAdvisor) Google (2022) was utilized to gather runtime metrics. This tool offers measurements related to resource usage, such as CPU, memory, disk Input/Output (I/O), and network I/O of Docker containers. Additionally, data from cAdvisor was also polled from the Prometheus server. Data collection spanned over six days under normal conditions. For each type of anomalous behaviour, data was collected over 90 minutes by injecting faults into specific services to simulate anomalies. Table 1 shows the features available in the RS-Anomic dataset. To ensure RS-Anomic covers an actual e-commerce application behaviour, we simulated a time-varying load with more users in peak hours and fewer in off-peak hours. We used Locust Locust (2022), an open-source load testing tool, for load generation. The load generation can be done by creating test scenarios that simulate user behaviour and defining the number of virtual users and requests per second that should be generated during the test.
Feature Description
container_memory_rss Resident set size of a container in bytes at the polled time
container_memory_usage_bytes Total memory usage in bytes at the polled time
container_memory_failures_total Cumulative count of memory allocation failures
container_memory_working_set_bytes Current working set size in bytes at the polled time
container_memory_failcnt Cumulative count of memory usage exceeds limit
container_cpu_usage_seconds_total Cumulative CPU usage seconds in total
container_cpu_user_seconds_total Cumulative user cpu time consumed
container_cpu_system_seconds_total Cumulative system cpu time consumed
container_network_receive_bytes_total Cumulative number of bytes received
container_network_receive_errors_total Cumulative count of errors encountered while receiving
container_network_receive_packets_dr-
opped_total
Cumulative count of packets dropped while receiving
container_network_receive_packets_total Cumulative count of packets received
container_network_transmit_bytes_total Cumulative number of bytes transmitted
container_network_transmit_errors_total Cumulative count of errors encountered while transmitting
container_network_transmit_packets_d-
ropped_total
Cumulative count of packets dropped while transmitting
container_network_transmit_packets_t-
otal
Cumulative count of packets transmitted
container_fs_usage_bytes Bytes used by a container on the file systems at the polled time
container_fs_io_time_seconds_total Cumulative time spend on file I/O
container_fs_write_seconds_total Cumulative time spent on file writes
Response times* Cumulative Response times to communicating services
Table 1: RS-Anomic feature descriptions. *Number of response times features may vary for each service based on the microservice architecture
Figure 1: RobotShop microservice architechture
3.1.2 Anomalous Behaviours Captured in RS-Anomic
  • Service Down: When a microservice is not working and cannot respond to requests, it is referred to as a service down. Network problems, hardware malfunctions, or software glitches can cause service down. To simulate this issue, we manually terminated the service and observed how the remaining services continued functioning.
  • High Concurrent User Load: When a large unexpected number of users try to access the application simultaneously, it can cause slow response times, timeouts, or service failures. To simulate a high concurrent user load, we used our load generation script to simulate 1500 concurrent users accessing the application.
  • High CPU usage: A microservice with a poorly optimized algorithm, inefficient code, or increased load can cause the entire application to underperform, resulting in high CPU usage. To simulate high CPU usage, we ran 100 parallel threads that calculate the 1000000th Fibonacci number.
  • High File I/O: This anomaly occurs when a microservice performs excessive file I/O operations leading to performance issues, such as slow response times and high CPU usage. To simulate high file I/O, we used a thread to continuously read from and write to a file.
  • Memory Leak: This fault occurs when a microservice fails to release memory that is no longer needed. Over time, this can increase memory usage and cause the service to crash or become unresponsive. Stress-ng ColinIanKing (2022) is used to continuously allocate memory without deallocating, which will simulate a memory leak.
  • Packet Loss: This anomaly occurs when packets of data sent between the microservice and other systems are lost due to network issues which result in slow response times, errors, or incomplete transactions. We use Traffic Control(tc) Linux (2022) to simulate network conditions where most packets are lost.
  • Response Time Delay: Response time delay occurs when a microservice takes longer than usual to respond to requests due to various factors, including high CPU usage, increased user load, network latency, or software bugs. We simulate this behaviour by adding a delay time to API service calls.
  • Out-of-Order Packets: Out-of-order packets are a common occurrence in microservices due to the distributed nature of the system and the use of asynchronous communication methods.
  • Low Bandwidth: A significant deviation from the expected bandwidth usage can be caused by inefficient communication protocols, lack of load balancing, network hardware limitations, or improper network configurations.
  • High Latency: Microservices rely heavily on network communication to interact with each other, and any increase in latency can cause delays in the response time of the microservices, leading to a degraded user experience.
We observed that network packet loss, out-of-order packets and high latency anomalies do not clearly distinguish between normal and anomalous data. Therefore, we increased the strength of these anomalies. Figure 2 shows t-SNE plots of comparison between two strengths of anomalous behaviour. The t-SNE plot for 1000ms latency shows clear distinctions between normal and anomalous behaviour for increased strength, whereas 200ms latency anomaly is more challenging to distinguish from normal data points.
(a) 200ms latency
(b) 1000ms latency
Figure 2: t-SNE plots for normal(blue) and anomalous data (orange) of varying strengths.
3.2 GAL-MAD Model
In recent work, we observed that unsupervised approaches have shown promising improvements due to the lack of anomalous data availability. Reproducing anomalies for various microservice applications is often impractical. Therefore, we propose an unsupervised approach, that leverages the normal behaviour performance metrics of services to train the model.
Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) is an autoencoder model Figure 3 illustrates the overall architecture of GAL-MAD. Figure 4 depicts the encoder component consisting of two GAT layers to extract latent information regarding the input feature dependencies and reduce the dimension along the feature axis. The GAT considers attention to neighboring services and features with the help of the adjacency matrix that represents the architecture of the microservice application, Figure 5 visualizes the attention computation for the input features. A many-to-one LSTM is used to reduce the dimension along the time axis. Each LSTM uses a sequence length of 24, translating to 2-minute time windows polled every 5 seconds. A smaller time window would improve responsiveness at the expense of accuracy.
Figure 3: GAL-MAD Model architecture
Figure 4: Encoder component of the GAL-MAD model
Figure 5: GAT layer of the GAL-MAD
The decoder mirrors the encoder’s structure, comprising a one-to-many LSTM followed by two GAT layers, to reconstruct along the time axis and feature axis respectively. Since GAT cannot output the required three-dimensional shape with the exact number of output features in the reconstruction, due to the incompatibility of the number of outputs and the number of attention heads, we employed a linear layer at all output nodes to adjust the feature axis to the desired size.
Normal data was used to train the model, which led to the expected higher MSE for anomalous data during testing. At the end of the reconstruction, a sigmoid function was used to get the final prediction. The model was trained using a batch size of 360 and a learning rate of 0.001 with a decay of 0.5 per epoch using the Adam optimizer.
3.3 Anomalous Service Localization
Anomalous service localization was conducted using SHAP on anomalous model detections. The results are depicted in Table 4. The process of locating the anomalous service is as follows,
  1. 1.
    SHAP values are calculated for each time step of a period
  2. 2.
    Anomalously detected time steps are filtered from the period
  3. 3.
    SHAP values are aggregated along the time axis
  4. 4.
    Absolute sum is computed over the resulting matrix along the feature axis
    1. (a)
      The SHAP values for the response time feature and container_fs_usage_bytes are weighted 4 times to improve results
  5. 5.
    Service is deduced by the index of the maximum value in the resulting array
  6. 6.
    The anomalous feature is deduced by the index of the highest value in the respective service array
4 Results
4.1 Experiments on RS-Anomic Dataset
We evaluated the RS-Anomic dataset using state-of-the-art anomaly detection models, GDN Deng and Hooi (2021), MAD-GAN Li et al. (2019), Kitsune Mirsky et al. (2018) and a custom transformer-based reconstruction model for reference Vaswani et al. (2017). GDN uses cosine similarity to create a graph structure, predicts future values using graph attention-based forecasting, and calculates the graph deviation score. MAD-GAN is designed to detect anomalies in time series data and uses an LSTM-RNN discriminator and a reconstruction-based method. Kitsune is an unsupervised Network Intrusion Detection System that can detect local network attacks without supervision. We categorized anomalies as the positive class and compared the models using Accuracy (A), Recall (R), and Specificity (S) metrics. The normal data was split into an 80:20 ratio for training and testing. Only normal data was used to train models based on respective papers. We created three testing scenarios with normal:anomalous behaviour ratios of 95:5, 90:10, and 60:40. For each testing scenario, the test data was mixed with anomalous data, according to the respective ratios. Anomalous samples were balanced across different anomaly types. The difference in the number of response time features for each microservice, due to varying numbers of links between microservices, was unified by summing all recorded response times for each service to create one response time feature. If a service experienced a network delay in any communication link, this information was preserved by summing the response times. In addition to response times, we introduced a moving average of response times for 5 minute and 30 minute windows. The input dimensions for all the models were 264 for each time step assembled by concatenating 22 features, including moving averages for 12 services. Moreover, standard scaling to scale our data based on normal data distribution. We considered a window size of 24 data points. We trained all models for 20 epochs with Adam optimizer and learning rates 0.001, 0.0005, 0.1 and 0.05 for GDN, MAD-GANs, Kitsune and Transformer, respectively. The results for GDN, MAD-GAN, Vanilla transformer, Kitsune and GAL-MAD models on the RS-Anomic dataset are displayed in Table 2. According to the results, especially the recall value, state-of-the-art anomaly detection models are not very successful in detecting anomalies in microservices. This indicates that RS-Anomic is a challenging dataset, and further research is necessary to develop new Machine Learning models for microservice anomaly detection.
The RS-Anomic dataset includes performance metrics of a microservice application under both normal and anomalous behaviour. This dataset enables anomaly detection in microservice applications and allows for the development of an anomaly classification model, as it contains ten different types of anomalies. However, it is essential to note that our application was deployed on a single server, therefore, the RS-Anomic dataset does not capture the impacts of a distributed deployment. Most microservice applications operate in distributed environments with auto-scaling, which introduces additional complexities that are not reflected in the RS-Anomic dataset. Table 2 presents the results of our empirical study, evaluating the existing anomaly detection models using the newly introduced RS-Anomic dataset demonstrates that there is still room to develop more accurate anomaly detection models. The suboptimal performance of current models may be attributed to their inability to account for the graph-like structure inherent in the RS-Anomic dataset. This structure, derived from the architecture of microservice applications, is critical for identifying dependencies among input features.
Model 95:5 90:10 60:40
A R S A R S A R S
GDN 0.9961 0.8093 0.9645 0.9409 0.7684 0.9620 0.8361 0.7766 0.8741
MAD-GANs 0.6073 0.8627 0.5920 0.5669 0.6893 0.5521 0.5225 0.5238 0.5258
Kitsune 0.9757 0.9006 0.9804 0.8984 0.5529 0.9412 0.5921 0.4142 0.7077
Transformer 0.9810 0.8252 0.9905 0.9678 0.7994 0.9885 0.8712 0.7491 0.9506
GAL-MAD 0.9838 0.9884 0.9559 0.9788 0.9228 0.9855 0.9500 0.8938 0.9859
Table 2: Performance of GAL-MAD and the state-of-the-art anomaly detection models on RS-Anomic dataset in terms of Accuracy(A), Recall(R), and Specificity(S) testing under a mixture of normal to anomaly ratios of 95:5, 90:10 and 60:40
4.2 Experiments on GAL-MAD
The GAL-MAD detector predicts anomalies in 80-second windows. Normal data windows produce a loss lower than 2.0. The reconstruction loss of the autoencoder model is passed through a sigmoid function to obtain the prediction as follows:
$prediction=sigmoid(loss-c)$
Where $loss$ is the reconstruction loss and $c$ is the upper bound of the reconstruction loss observed for normal data. Table 3 shows average loss values for two configurations of the model, one using only cAdvisor metrics and the other using both cAdvisor metrics and response time features. The response time features include moving averages calculated over 5 minute and 30 minute windows. The analysis reveals the significant impact of incorporating response time features on anomaly detection, particularly for network-related anomalies. Moreover, packet loss, out-of-order packets, and low bandwidth anomalies were tested under two different strengths to highlight the lack of sensitivity observed on fine network anomalies.
Anomaly/Normal Model without response times Model with response times
Normal 1.974 1.930
Service Down 17554.017 40366.418
High CPU usage 353297.431 338129.254
High Concurrent User
Load (1500 users)
30781.602 31068.059
High File I/O 278496.914 264153.407
Memory Leak (upto
300mb)
242.596 224.751
Packet Loss(50%) 2.045 157.540
Packet Loss(80%) 11.597 23703.363
Response Time
Delay($\sim$400ms)
1.854 27714.654
Out of Order
Packets(25%)
2.080 2.489
Out of Order
Packets(60%)
1.677 23694.171
Low Bandwidth(1kbps
burst 256b)
7.102 10.704
Low Bandwidth (1kbps
burst 64b)
9.869e+16 9.272e+16
High Latency(200ms) 1.984 29.678
High latency(1200ms) 1.440 23693.894
Table 3: Losses with and without response time for each anomaly type
4.2.1 Comparative Study of GAL-MAD Model
We conducted a comparative study with the state of the art anomaly detection models on our dataset. The results in Table 2 show that our model has better performance, given the complex nature of the data.
4.3 Anomalous Service Localization Scores
Anomaly localizations are considered true when both the service with the injected anomaly and its related features are successfully identified. SHAP values visualized as heatmaps for selected anomaly instances are shown in Figures 6 to 9. The features are indicated by labels 0 to 21 in the X axis, listed in order as shown in Table 1, followed by the moving averages of response times in 5 minute and 30 minute windows labeled as 20 and 21, respectively. The Y axis represents the payment, shipping, redis, mongodb, dispatch, rabbitmq, user, mysql, catalog, ratings, web and cart services in order.
The heatmaps provide intuitive explanations in cases related to performance and response time anomalies. Figure 6 illustrates a heatmap with diverging shap values for response time features when the API calls from the catalogue service to the mongodb service were delayed. Figure 7 demonstrates that CPU usage related features of the dispatch service with a diverging SHAP value when an abnormally high CPU usage is introduced.
However, it is notable that fine network-related anomalies such as out-of-order-packets and packet loss are substantially more difficult to deduce from visualizations Figures 8 and 9. The SHAP values were modified by increasing the weight of container_fs and response times features fourfold, with results documented in Table 4.
Anomaly True service localization True feature localization
rt-delay 10/10 10/10
high-cpu 10/10 10/10
high-fileIO 10/10 10/10
memory-leak 10/10 3/10
low-bandwidth 0 0
out-of-order-packets 0 0
high-latency 6/10 6/10
packetloss 9/10 9/10
Table 4: Anomalous service and feature localization results
5 Conclusion
This research introduces the RS-Anomic dataset, a comprehensive multivariate time-series resource designed to advance anomaly detection research in microservices architectures. The dataset encompasses 100,000 normal data points and 14,000 anomalous data points, spanning ten distinct anomaly types, providing a robust foundation for academic and industrial researchers. Furthermore, we propose the GAL-MAD model for anomaly detection, which leverages Graph Attention Networks and Long Short-Term Memory networks to capture both spatial and temporal dependencies between microservices. Our empirical study evaluated existing anomaly detection models using the RS-Anomic dataset. The GAL-MAD model outperformed the state-of-the-art anomaly detection on the RS-Anomic dataset. The results of using SHAP for explainability show that it facilitates localizing the anomalous service and the feature to a great extent in performance-related anomalies. System administrators can leverage such a framework to quickly identify and diagnose anomalies, understand the root cause of performance degradations and, gain deeper insight to microservice interactions. The RS-Anomic dataset and GAL-MAD model represent important steps toward more intelligent and interpretable anomaly detection in complex distributed systems. Further research directions include expanding the dataset to incorporate multi-node architectures and auto-scaling capabilities that more closely reflect production-level microservices deployments. Additionally, while SHAP provides insights into root cause localization, further work is required in interpreting anomalies by integrating the application architecture context. Further analysis of network-related metrics is required to improve network anomaly explainability and diagnostic precision.
Figure 6: Heat map for response time delay anomaly in the mongo database service. Service 3 represents the MongoDB service. Features 19, 20, and 21 represent response times and moving averages of the response times for 5 and 30 minute windows, respectively.
Figure 7: Heat map for high CPU usage in the dispatch service anomaly. Service 4 represents the dispatch service. Features 1,2,3, and 5 represent memory usage, memory failures, memory working set usage and CPU usage time, respectively.
Figure 8: Heat map for a false localization of out-of-order packets in the user service.
Figure 9: Heat map for a false localization of packet loss anomaly in the user service.